Calculate the initial molar concentration of HF solution with a pH of 2.21.
(Kb = 1.5 x 10à ƒ ƒ ƒ ¢ € ’11)
1
Expert's answer
2021-10-15T02:55:16-0400

2.21= 14-Log(1.5× 10)X

2.21= 14-1.6(X)

1.6X = 14-2.21

1.6X= 11.79

= 7.37mol/L